# Dynamic and Distributed Routing in IoT Networks based on Multi-Objective Q-Learning

### v2

@@ VII Experimental Results
We compare the efficiency of the proposed Distributed DPQ-Learning algorithm with six baseline schemes listed below:
@@
    Reinforcement-Learning-Based Energy Efficient Control and Routing Protocol for Wireless Sensor Networks (RLBEEP) [abadi2022rlbeep]
@@
    Reinforcement Learning for Life Time Optimization (R2LTO) [bouzid2020efficient]
@@
    Reinforcement Learning Based Routing (RLBR) [guo2019optimizing]
@@ VII-A The Simulation Setup and Parameters
The variation of $\epsilon$-values and preference variation are shown in Figure 2. Since experiments 1 and 2 are characterized by sequential exploration and exploitation, the $\epsilon$ value (Figure 2(a)) in the proposed algorithm is $1$ in the first $1000$ episodes (indicating full exploration and no exploitation) and 0 afterward (indicating no exploration and only exploitation). In experiments $3$ and $4$ (Figure 2(b)), the $\epsilon$ is set as $1$ in the beginning and then linearly reduced to 0, thus achieving the desired simultaneous exploration and exploitation. In all the experiments, the $\epsilon$ values for the SMORLR baseline are reset to $1$ each time the decision-maker changes the preference. Then, it linearly reduces to 0, encouraging fresh exploration after each preference-change. The frequency of preference variation is a thousand episodes in experiments $1$ and $3$ (Figure 2(c)), and each episode in experiments $2$ and $4$ (Figure 2(d)). The two different preference variation patterns are considered to study the effect of the frequency of preference changes on performance.
@@ VII-B2 Energy
The cumulative effects over the long run can be more clearly seen in Figure 4(b), which depicts the cumulative energy consumption over episodes. The figure shows that the proposed algorithm is the most energy-efficient, consuming around $4\times$ and $3.4\times$ less than RSPAR and FMOLD respectively, $30-50\%$ lesser energy than SMORLR and RLBEEP, and $3-5\%$ lesser than RLBR, and R2LTO by episode $10000$. While SMORLR, RLBEEP, R2LTO, and RLBR perform moderately well, RSPAR and FMOLD exhibit the highest cumulative energy consumption, highlighting their inefficiency. However, the proposed routing algorithm still consumes less energy than any baseline algorithm. Similarly we continue our evaluations in Experiment 2, where preferences were randomly changed in each episode. As shown in Figure 4(c), the cumulative energy consumption for the proposed routing protocol is lesser than those of the baseline methods, with the gap widening as more episodes progress. By the end of 100 episodes, we find that the cumulative energy consumption of RSPAR and RLBEEP is $3.87\times$, FMOLD is $3.53\times$, RLBR is $2\times$, R2LTO is $1.73\times$, and SMORLR is $1.67\times$ that of the proposed routing scheme. Therefore, regardless of the exploration policy or preference variation patterns, the proposed method leads to a substantially lesser amount of energy consumption than the baseline algorithms.
@@ VII-C1 Overall Reward
We present both episodic rewards (figure 6(a)) and accumulated returns (figure 6(b)) of the proposed Distributed DPQ-Learning routing algorithm against the six baseline routing algorithms. From figure 6(a), we can observe that the proposed Distributed DPQ-Learning routing algorithm learns to optimize the overall return by doing more exploration than exploitation in the first $1000$ episodes. The pareto front is also learned in this phase. In contrast to the sequential exploration-exploitation scheme, the simultaneous scheme involves both exploration and exploration occurring hand-in-hand while gradually reducing exploration and increasing exploitation. From figure 6(a), we notice the performance of the proposed Distributed DPQ-Learning routing algorithm may not be superior in the beginning phases of exploration, but after a few exploration ($1000$ episodes), it outperforms all the baselines with respect to episodic rewards. The proposed protocol also dynamically and immediately adjusts to changing preferences after a few explorations, while other baselines fail to do so. For example, SMORLR learns to adapt after each preference change. However, the process is slow, and the preferences change before it can adapt to match the proposed protocol’s performance. Further, R2LTO, RLBEEP, FMOLD, and RLBR consistently focus on a fixed objective preference, mostly prioritizing energy efficiency. They tend to learn certain intervals to increase returns, but they still fall short of the proposed method by $12-13\times$ approximately. RSPAR performs, on average, 344% less efficiently than the proposed approach. Thus, while certain baseline methods may perform well under specific preferences, the proposed algorithm consistently outperforms them and adapts quickly to changing preferences.
@@ VII-C3 Packet Delivery
Figure 8(b) provides a more comprehensive view by showcasing cumulative packet delivery, which offers a fair comparison over the long run. The proposed method starts outperforming all baselines, after $3000$ episodes of sufficient learning. As time progresses, the performance gap widens. By the end of $10,000$ episodes, the proposed method delivers $46.67\%$ more packets than RSPAR, $69.23\%$ more packets than SMORLR, $164\%$ more packets than FMOLD, $230\%$ more packets than RLBR, and $255.12\%$ more packets than R2LTO and RLBEEP. Similar results were seen in Experiment 4, where preferences were randomly changed in each episode. As depicted in Figure 8(c), the total number of packets delivered by the proposed routing protocol consistently and significantly exceeds those of the baselines, with the gap widening as more episodes are completed. Specifically, at the end of $5000$ episodes, the total packet delivery for the proposed algorithm reaches around $3200$, which is $33.33\%$ more than RSPAR, around $166\%$ more than FMOLD and RLBR, $220\%$ more than SMORLR, and $814.3\%$ more than more than RLBEEP and R2LTO. The proposed method outperforms all baselines when the preference for PDR objective is higher, and is also acceptable for preferences at lower levels.
@@ VIII Conclusions
This paper presents a distributed multi-objective Q-learning–based routing algorithm for IoT and wireless sensor networks. The method learns multiple per-preference Q-tables in parallel and uses a greedy interpolation policy to adapt instantly to changing preference weights without retraining or central coordination. Theoretical analysis establishes a uniform near-optimality guarantee, explaining the stability and effectiveness of the proposed interpolation mechanism. Simulation results confirm consistent gains over existing schemes, demonstrating that the proposed Dynamic Preference Q-Learning framework provides a light-weight and theoretically grounded approach for real-time distributed routing in dynamic IoT environments.
Shubham Vaishnav is currently pursuing Ph.D. at the Department of Computer and Systems Science, Stockholm University, Sweden. This research program is funded by Digital Futures. He received Bachelors’ and Masters’ degrees in Computer Science and Engineering from the Indian Institute of Technology (Indian School of Mines), Dhanbad. His research interests include Multiobjective and Adaptive decision-making in IoT and Edge networks, Reinforcement learning, Federated Learning, and data-driven decision-making in critical societal infrastructure.
Praveen Kumar Donta (SM’22), Associate Professor (Docent) at the Department of Computer and Systems Sciences, Stockholm University, Sweden. Earlier, he was a Postdoc at Distributed Systems Group, TU Wien, Vienna, Austria. He received his Ph. D. at the Indian Institute of Technology (Indian School of Mines), Dhanbad, from the Department of Computer Science and Engineering in May 2021. He was a visiting PhD student at the University of Tartu, Estonia, during his PhD for six months. He received his Bachelor’s and Master’s in Technology from JNT University, Anantapur, India. He is an Editorial Board member for IEEE IoT Journal, Nature Scientific Reports, Computing Springer, and ETT Wiley. His current research includes the Learning techniques for Distributed Computing Continuum Systems.
Sindri Magnússon is an Associate Professor in the Department of Computer and Systems Science at Stockholm University, Sweden. He received a B.Sc. degree in Mathematics from the University of Iceland, Reykjav´ık Iceland, in 2011, a Masters degree in Applied Mathematics (Optimization and Systems Theory) from KTH Royal Institute of Technology, Stockholm, Sweden, in 2013, and the PhD in Electrical Engineering from the same institution, in 2017. He was a postdoctoral researcher 2018-2019 at Harvard University, Cambridge, MA, and a visiting PhD student at Harvard University in 2015 and 2016. His research interests include large-scale distributed/parallel optimization, machine learning, and control.

### v4

@@ VII Experimental Results
We compare the efficiency of the proposed Distributed DPQ-Learning algorithm with the following baselines:
@@
    Reinforcement-Learning-Based Energy Efficient Control and Routing Protocol for Wireless Sensor Networks (RLBEEP) [1]
@@
    Reinforcement Learning for Life Time Optimization (R2LTO) [3]
@@
    Reinforcement Learning Based Routing (RLBR) [12]
@@ VII-A The Simulation Setup and Parameters
The experiments are conducted using a custom Python-based simulator (Python 3.9.13) executed on an 11th Gen Intel(R) Core(TM) i7-1185G7 CPU @ 3.00GHz with 32 GB RAM on Windows. The simulated network follows an IEEE 802.15.4 MAC abstraction over a $10\times 10$ grid topology ($100$ nodes), where each grid cell corresponds to an IoT node and edges represent one-hop connectivity. This corresponds to a unit-disk communication model where each node can directly communicate with its four adjacent neighbors (one-hop radio range). With the adopted local DPQ representation, the simulated per-node routing memory is approximately $1.72$ KB for the fine preference grid ($|\mathcal{B}|=11$) and $320$ bytes for the coarse grid ($|\mathcal{B}|=2$), assuming 64-bit floating-point Q-values.
@@
We evaluate four settings defined by two orthogonal factors: (i) exploration strategy and (ii) preference-change frequency, as summarized in Fig. 2. For exploration, Experiments 1–2 use sequential exploration/exploitation, where the proposed algorithm applies $\epsilon=1$ for the first $1000$ episodes (exploration) and $\epsilon=0$ thereafter (exploitation). Experiments 3–4 use simultaneous exploration/exploitation, where $\epsilon$ is initialized at $1$ and linearly decayed to 0 over time. For preference variation, Experiments 1 and 3 update the preference periodically every $1000^{\text{th}}$ episode (periodically changing objectives), while Experiments 2 and 4 update the preference at every episode (rapidly changing objectives). For the SMORLR baseline, $\epsilon$ is reset to $1$ after each preference update and then linearly decayed to 0, enabling fresh exploration under the new objective.
@@ VII-B2 Energy
[…]
@@ VII-C1 Overall Reward
We present both episodic rewards (Figure 6(a)) and accumulated returns (Figure 6(b)) of the proposed Distributed DPQ-Learning routing algorithm against the six baseline routing algorithms. From Figure 6(a), we observe that the proposed algorithm learns to optimize the overall return by performing more exploration than exploitation in the first $1000$ episodes, during which the Pareto front is also learned. In contrast, the simultaneous exploration–exploitation scheme carries out both processes in parallel, while gradually reducing exploration and increasing exploitation.
[…]
@@ VII-C3 Packet Delivery
Figure 8(b) provides a more comprehensive view through cumulative packet delivery, enabling a fair long-run comparison. The proposed method begins to outperform all baselines after $\approx 3000$ episodes of sufficient learning, and the gap widens over time. By the end of 10,00010{,}000 episodes, it delivers $46.67\%$ more packets than RSPAR, $69.23\%$ more than SMORLR, $164\%$ more than FMOLD, $230\%$ more than RLBR, and $255.12\%$ more than R2LTO and RLBEEP. Similar trends appear in Experiment 4 with per-episode random preference changes. As shown in Figure 8(c), the proposed protocol consistently exceeds all baselines, with the separation increasing as episodes progress. At $5000$ episodes, it reaches around $3200$ delivered packets, which is $33.33\%$ higher than RSPAR, around $166\%$ higher than FMOLD and RLBR, $220\%$ higher than SMORLR, and $814.3\%$ higher than RLBEEP and R2LTO. Overall, the proposed method performs best when the preference for PDR is high, while remaining acceptable for lower preferences.
VII-D Sensitivity Analysis
To assess the robustness of the proposed Dynamic Preference-based Q-learning (DPQ) routing framework, we conducted a sensitivity analysis by varying the temporal aggregation window used during testing. This analysis evaluates how performance changes when network preferences persist for shorter or longer durations, which is representative of heterogeneous and time-varying IoT operating conditions.
Experimental Setup
All experiments were conducted in the same wireless sensor network environment described earlier. After a fixed training phase, routing policies were evaluated over testing windows of length $W\in\{50,200,500\}$ episodes. To evaluate sensitivity with respect to preference discretization, the proposed method was tested using two different preference grids: a coarse grid with two anchor points, $\beta\in\{0,1\}$, and a fine grid with eleven anchor points, $\beta\in\{0,0.1,\ldots,1\}$. All other parameters, including network topology, traffic patterns, and random seeds, were kept identical across methods to ensure a fair comparison.
The cumulative reward, which jointly captures packet delivery success and energy consumption, is the primary optimization objective of the proposed method. Accordingly, reward is reported first and highlighted in the tables, while packet delivery ratio (PDR) and energy consumption are included as supporting metrics.
Compared Methods
We compare two variants of the proposed approach—DPQ with a coarse preference grid (DPQ coarse_2) and a finer grid (DPQ fine_11)—against representative reinforcement-learning-based baselines (R2LTO, RLBR, RLBEEP) and a static multi-objective Q-learning method (SMORLR). The coarse and fine grids correspond to two-point and eleven-point preference discretizations, respectively.
Results and Discussion
Tables IIIV summarize the sensitivity results. Across all window lengths, the proposed DPQ variants consistently achieve the highest cumulative reward, demonstrating robustness to changes in preference persistence. While certain baselines occasionally achieve comparable PDR or lower energy consumption, they fail to balance these objectives effectively, resulting in significantly lower reward. This confirms that optimizing individual metrics in isolation does not yield optimal long-term routing performance.
TABLE II: Sensitivity results for window length $W=50$ (mean $\pm$ std).
Method Reward ($\uparrow$) PDR ($\uparrow$) Energy ($\downarrow$)
DPQ coarse_2 0.863 $\pm$ 0.236 0.954 $\pm$ 0.168 0.196 $\pm$ 0.331
DPQ fine_11 0.862 $\pm$ 0.236 0.953 $\pm$ 0.161 0.110 $\pm$ 0.011
SMORLR (Static-Q) 0.120 $\pm$ 0.066 0.141 $\pm$ 0.056 0.193 $\pm$ 0.026
R2LTO 0.120 $\pm$ 0.063 0.142 $\pm$ 0.052 0.182 $\pm$ 0.020
RLBR 0.120 $\pm$ 0.063 0.142 $\pm$ 0.052 0.182 $\pm$ 0.020
RLBEEP 0.120 $\pm$ 0.063 0.142 $\pm$ 0.052 0.182 $\pm$ 0.020
TABLE III: Sensitivity results for window length $W=200$ (mean $\pm$ std).
Method Reward ($\uparrow$) PDR ($\uparrow$) Energy ($\downarrow$)
DPQ coarse_2 0.610 $\pm$ 0.374 0.815 $\pm$ 0.298 0.437 $\pm$ 0.579
DPQ fine_11 0.605 $\pm$ 0.374 0.809 $\pm$ 0.285 0.101 $\pm$ 0.017
SMORLR (Static-Q) 0.292 $\pm$ 0.223 0.403 $\pm$ 0.158 0.140 $\pm$ 0.025
R2LTO 0.350 $\pm$ 0.185 0.545 $\pm$ 0.031 0.150 $\pm$ 0.005
RLBR 0.350 $\pm$ 0.185 0.545 $\pm$ 0.031 0.150 $\pm$ 0.005
RLBEEP 0.350 $\pm$ 0.185 0.545 $\pm$ 0.031 0.150 $\pm$ 0.005
TABLE IV: Sensitivity results for window length $W=500$ (mean $\pm$ std).
Method Reward ($\uparrow$) PDR ($\uparrow$) Energy ($\downarrow$)
DPQ coarse_2 0.453 $\pm$ 0.346 0.737 $\pm$ 0.346 0.516 $\pm$ 0.653
DPQ fine_11 0.439 $\pm$ 0.346 0.713 $\pm$ 0.332 0.095 $\pm$ 0.019
SMORLR (Static-Q) 0.232 $\pm$ 0.227 0.417 $\pm$ 0.212 0.101 $\pm$ 0.013
R2LTO 0.413 $\pm$ 0.246 0.819 $\pm$ 0.011 0.129 $\pm$ 0.003
RLBR 0.413 $\pm$ 0.246 0.819 $\pm$ 0.011 0.129 $\pm$ 0.003
RLBEEP 0.413 $\pm$ 0.246 0.819 $\pm$ 0.011 0.129 $\pm$ 0.003
@@ VIII Conclusions
This paper presents a distributed multi-objective Q-learning–based routing algorithm for IoT and wireless sensor networks. The method learns multiple per-preference Q-tables in parallel and uses a greedy interpolation policy to adapt instantly to changing preference weights without retraining or central coordination. Theoretical analysis establishes a uniform near-optimality guarantee, explaining the stability and effectiveness of the proposed interpolation mechanism. Simulation results confirm consistent gains over existing schemes, demonstrating that the proposed Dynamic Preference Q-Learning framework provides a light-weight and theoretically grounded approach for real-time distributed IoT routing.The sensitivity analysis validates that the proposed method maintains superior reward performance across diverse settings, highlighting its robustness and suitability for dynamic IoT routing scenarios with time-varying objectives.